OpenStudy (32561):

use the following function rule to find f(39) f(x)=15-x/13

OpenStudy (mathmale):

Start with f(x)=15-x/13. Replace each "x" with "39." Show your result. Then simplify your result.

OpenStudy (grace00now):

f(39)=12

OpenStudy (32561):

how if f=39

OpenStudy (32561):

f(48) f(x)=9-x/12

OpenStudy (grace00now):

f(48)=5

OpenStudy (grace00now):

This is super easy :)

OpenStudy (32561):

f(50) f(x)=11-x/10

OpenStudy (32561):

f(r=8) f(n)=n+12